OpenStudy (anonymous):

(x^2-9)/(2x^2+7x+3) * (2x^2+11x+5)/(x^2-3x)

OpenStudy (anonymous):

is this 2 problems

OpenStudy (anonymous):

No it is all one problem. The left one multiplied by the right one.

OpenStudy (anonymous):

oh sorry I see the * is multiply

OpenStudy (skullpatrol):

$$\huge \frac{x^2-9}{2x^2+7x+3} * \frac{2x^2+11x+5}{x^2-3x}$$

OpenStudy (anonymous):

Yea.

OpenStudy (skullpatrol):

Looks like you have some factoring to do :)

OpenStudy (anonymous):

x^2-9 is factored into (x+3) (x-3)

OpenStudy (anonymous):

I'm lost after this?

OpenStudy (skullpatrol):

factor an x out of $$x^2 -3x$$

OpenStudy (anonymous):

(x-3)?

OpenStudy (anonymous):

somebody?

OpenStudy (agent0smith):

so far...\[\huge \frac{(x+3)(x-3)}{2x^2+7x+3} * \frac{2x^2+11x+5}{x(x-3)}\]

OpenStudy (anonymous):

How do I factor 2x^2+7x+3?

OpenStudy (agent0smith):

2x^2+7x+3 multiply the first and last numbers: 2*3 = 6 now find two numbers that multiply to equal 6, and add up to the middle number (7).

OpenStudy (anonymous):

6*1. So (x+6) (x+1)?

OpenStudy (agent0smith):

It's not that simple this time. break up the 7x like this: \[\Large 2x^2+7x+3\] \[\Large 2x^2+6x+1x+3\] \[\Large 2x(x+3)+1(x+3)\]

OpenStudy (anonymous):

Ok, so how do I break up the 10 and one for the next part?

OpenStudy (agent0smith):

Do you know how to finish factoring this \[\Large 2x(x+3)+1(x+3)\]

OpenStudy (anonymous):

x(x+3)?

OpenStudy (agent0smith):

Take out the common factor (x+3), what gets left behind? 2x(x+3)+1(x+3) (x+3)(2x+1)

OpenStudy (anonymous):

Ok, I got that.

OpenStudy (agent0smith):

\[\huge \frac{(x+3)(x-3)}{(x+3)(2x+1)} * \frac{2x^2+11x+5}{x(x-3)}\] try factoring the 2x^2+11x+5 using a similar method to above.